OpenStudy (anonymous):

Choose the most appropriate translation of the English sentence. The circumference of a balloon (which depends on the amount of air inside) is 1 meter. A. Circumference(air) = 1 B. Circumference(air) < 1 C. Circumference(1) > 0 D. Air(circumference) = 1

OpenStudy (anonymous):

5x + y = 2

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

5x + y = 2 x-intercept y-intercept

OpenStudy (anonymous):

5x + y = 2 x-intercept y-intercept

OpenStudy (anonymous):

so is it like b or what ?